A stock has a beta of 1.6 and an expected return of 16%. A risk-free asset...

  1. A stock has a beta of 1.6 and an expected return of 16%. A risk-free asset currently earns 5%.
  1. If a portfolio of the two assets has a beta (βp) of 0.6, what are the portfolio weights? (wS = 37.5%, wRF = 62.5%)
  2. If a portfolio of the two assets has an expected return of 11%, what is its beta? (βp= 0.873)
Answer #1

The solution to the given question is provided below:

1.

w*1.6+(1-w)*0=0.6

=>w=0.6/1.6

w1=37.5%

w2=1-w1=62.5%

2.

=(11%-5%)/(16%-5%)/1.6

=0.873